How many photons per second are emitted by a 5.5-W LED light bulb, if the emission is at a wavelength of 550 nm?

~ 1.23E22 photons/s

~ 5.07E21 photons/s

~ 3.62E19 photons/s

~ 1.52E19 photons/s

~ 4.27E19 photons/s
Answer #1

energy of each photon

E = hc/L

now

Number of photons * energy in each / time = Power

N* E / t = P

(N/t) * hc/L = P

(N/t) * 6.626* 10^-34* 3* 10^8 / ( 550* 10^-9) = 5.5

N/t = 1.52* 10^19 photons /s

=========

Comment before rate in case any doubt, will reply for sure.. goodluck